# Bridging the Culture Gap: A Framework for LLM-Driven Socio-Cultural Localization of Math Word Problems in Low-Resource Languages

### v2

@@ 6 Conclusion
Acknowledgment
The authors would like to thank the German Federal Ministry of Education and Research and the German federal states (http://www.nhr-verein.de/en/our-partners) for supporting this work/project as part of the National High-Performance Computing (NHR) joint funding program. We also thank PaliGemma academic program for gemini credits used in this research.
@@ Acknowledgment
Appendix A Prompt Templates Used
@@
Prompt: Entity Localization Task You are an expert linguistic assistant. Your task is to edit a sentence in a native language to match a change made in its English parallel.
Here is the context: Original English: The original sentence. Original Native: The original translation of the English sentence in {native_lang}. Modified English: The English sentence has been edited. One or more words have been replaced. Your goal is to produce a Modified Native sentence by applying the exact same replacement to the Original Native sentence.
@@
Crucial Instructions: DO NOT re-translate the entire sentence. Only replace the specific words that were changed in the English version. Preserve the original grammar and structure of the native sentence as much as possible. Ensure the final Modified Native sentence is natural and grammatically correct in {native_lang}. Respond with ONLY the Modified Native sentence and nothing else.
Example: Original English: Janet’s ducks lay 16 eggs per day. Original Native (French): Les canards de Janet pondent 16 œufs par jour. Modified English: Andrea’s ducks lay 16 eggs per day. Modified Native (French): Les canards d’Andrea pondent 16 œufs par jour.
Your Task: Original English: {original_eng} Original Native ({native_lang}): {original_native} Modified English: {modified_eng} Modified Native ({native_lang}):
Prompt: Direct Localization using LLMs without our pipeline You are an expert linguistic assistant specializing in localization. Your task is to localize the following English sentence into {target_lang_name}.
@@
Localization means adapting the text to the target culture and context by: Identifying and replacing English-specific entities (such as person names, organization names, and currency symbols/names) with culturally appropriate equivalents in {target_lang_name}. Ensuring the localized text sounds natural and grammatically correct in {target_lang_name}. Crucially, the localized text should be generated directly from the English sentence, without relying on any pre-existing native translation. English Sentence to Localize: {original_eng} Your Goal: Produce a Localized {target_lang_name} sentence. Instructions: Identify person names, organization names, and currency symbols/names in the English sentence. Replace these entities with culturally appropriate {target_lang_name} alternatives. If an entity does not have a direct cultural equivalent or is already culturally neutral, it may remain unchanged. Respond with ONLY the Localized {target_lang_name} sentence and nothing else. Localized {target_lang_name}:
Appendix B Additional Evaluation Results
Figure 6 presents additional evaluations, extending the analysis to a broader set of models with diverse characteristics.
Figure 6: Translated vs Auto Localization performance Gemma-2-2b-it (top left) Llama-3.2-1B-Instruct (top right) Qwen3-4B (bottom left) LLaMA-3-8B-Instruct (bottom right)
Appendix C Precence of Cultural entities in our Training data
@@
Figure 7 illustrates the distribution of culturally relevant items within the 1,500 data samples selected for each language. Our analysis reveals that a considerable proportion of the translated items could not be localized to the target language, primarily because they did not contain the types of entities, such as personal names, organizations, or currencies, that form the basis of our localization strategy. This highlights a key limitation of translation-only benchmarks: when source items lack culturally grounded entities, they fail to provide meaningful opportunities for localization, thereby reducing the effectiveness of culturally adapted datasets. By quantifying this gap, we emphasize the importance of designing datasets that explicitly incorporate entity-level cultural relevance to ensure robust evaluation and training across languages.
@@
Figure 7: Number of Data Samples Without Cultural Entity Replacements out of 1500 selected, Grouped by Language for Training dataset.
Appendix D Translation quality - SSA-COMMET score
@@
Appendix E Code and Reproducibility

### v3

@@ 6 Conclusion
Appendix A Full Sampling result
As shown in Table 4, the results provide a comprehensive view of how different sampling sizes and training configurations affect robustness across languages. Several key observations emerge. First, the effect of data localization is not uniform across models: while LLaMA-3-8B-Instruct often exhibits fluctuations depending on sample size, Gemma-2-9b-it tends to benefit more consistently from auto-localized data ($\hat{x}_{\text{loc}}$). This suggests that Gemma is more sensitive to entity grounding and gains robustness when trained with culturally and linguistically aligned examples.
Lang LLaMA-3-8B-Instruct Gemma-2-9b-it
$x_{\text{trans}}$ $x_{\text{ent}}$ $\hat{x}_{\text{loc}}$ all data $x_{\text{trans}}$ $x_{\text{ent}}$ $\hat{x}_{\text{loc}}$ all data # sample
amh 0.13 -1.20 1.07 0.93 0.53 -0.53 -0.93 0.80 1000
amh -0.27 -0.13 1.33 0.40 0.27 0.00 0.40 -1.20 5000
amh 1.33 -0.27 0.53 -1.07 -0.13 0.40 -0.53 -0.13 10000
amh 1.33 -0.27 0.53 -1.07 -1.73 0.93 1.07 0.40 25100
ewe -0.67 0.13 0.27 0.80 0.27 -0.53 0.27 1.20 25100
ewe -0.67 0.13 0.27 0.80 0.27 -0.53 0.27 1.20 10000
ewe 0.40 0.27 -0.13 -0.67 0.13 0.13 0.27 -0.67 1000
ewe -0.27 -0.40 0.13 -1.33 0.53 -0.93 1.87 -0.67 5000
fra 1.07 0.93 0.67 -0.93 0.40 -0.53 -0.27 -0.13 5000
fra -0.53 0.67 -0.67 -0.67 -1.07 -1.33 -0.93 0.53 10000
fra -1.60 -0.53 -0.67 -0.67 -0.67 2.40 0.80 -0.93 1000
fra -0.53 0.67 -0.67 -0.67 -0.13 0.40 -0.53 -0.13 25100
hau -0.13 -0.27 -1.20 0.80 -2.13 -1.07 -1.87 0.00 5000
hau -0.80 -0.80 -1.20 -0.27 -1.73 0.93 1.07 0.40 10000
hau -0.93 -0.27 -1.60 -0.93 -0.40 -1.33 -0.67 -2.53 1000
hau -0.80 -0.80 -1.20 -0.27 -1.20 -0.93 0.67 -0.93 25100
ibo 0.00 -0.27 0.13 0.67 -0.53 1.87 0.67 0.27 1000
ibo 0.67 0.13 0.80 -0.80 -0.67 0.67 0.67 0.13 5000
ibo 0.53 0.80 0.67 -0.93 -0.93 0.67 0.27 -1.33 10000
ibo 0.53 0.80 0.67 -0.93 -0.40 -0.13 -0.40 -1.07 25100
kin 0.27 -0.27 -0.40 1.73 0.80 -0.13 1.33 3.60 1000
kin -1.07 -0.67 0.40 -0.27 0.00 -1.60 0.67 0.67 25100
kin 0.00 1.20 -0.53 0.53 -0.27 1.07 -1.73 -1.47 5000
kin -1.07 -0.67 0.40 -0.27 0.00 -0.13 -0.53 0.27 10000
lin -0.13 -0.67 0.27 0.67 -0.27 -0.13 -1.20 0.67 1000
lin -1.07 -0.80 -0.13 -0.67 0.40 0.67 2.00 1.47 10000
lin -1.07 -0.80 -0.13 -0.67 0.40 0.67 2.00 1.47 25100
lin -0.27 -0.40 -0.67 0.40 -1.60 0.53 0.80 0.13 5000
lug -0.13 0.13 0.00 0.40 -1.07 -1.47 -0.13 -0.13 1000
lug -0.80 -1.07 0.00 -1.07 0.13 -1.59 1.07 0.27 5000
lug -1.33 -0.13 -0.27 -1.33 -2.40 -0.26 0.93 -1.33 10000
lug -1.33 -0.13 -0.27 -1.33 -2.40 -0.26 0.93 -1.33 25100
orm -0.40 -0.80 -0.53 0.67 0.40 -0.40 -0.13 -1.07 10000
orm -0.40 -0.80 -0.53 0.67 0.40 -0.40 -0.13 -1.07 25100
orm -0.40 0.27 -0.80 -0.53 -0.27 -0.40 0.80 -1.07 1000
orm -0.93 -0.53 -1.33 -0.27 0.27 0.80 -0.27 -1.07 5000
sna -1.73 -0.27 -0.13 0.93 2.27 -0.27 1.20 1.60 1000
sna 0.00 1.07 0.13 -0.80 0.80 0.13 1.87 -0.40 5000
sna -2.00 -0.53 -0.53 0.13 -0.13 0.53 -0.40 0.27 10000
sna -2.00 -0.53 -0.53 0.13 0.00 -0.13 -0.53 0.27 25100
sot -0.40 0.00 0.40 0.53 1.07 0.80 2.93 -0.13 1000
sot 0.40 0.80 -1.33 0.27 0.00 -1.60 0.67 0.67 10000
sot 0.00 0.67 -0.80 1.60 0.67 -1.07 0.00 -1.33 5000
sot 0.40 0.80 -1.33 0.27 -1.07 -1.33 -0.93 0.53 25100
swa 0.40 1.33 -0.53 0.67 -0.40 0.00 0.27 -0.13 1000
swa -0.67 -1.60 -0.13 0.67 -2.00 1.60 1.20 1.07 5000
swa -0.67 -1.20 0.40 0.27 -1.07 -1.33 0.40 -0.40 10000
swa -0.67 -1.20 0.40 0.27 -1.07 -1.33 0.40 -0.40 25100
twi 0.67 -0.27 0.53 -0.93 2.00 0.53 1.47 -0.40 10000
twi 0.67 -0.27 0.53 -0.93 2.00 0.53 1.47 -0.40 25100
twi 0.53 -1.07 -0.13 -0.40 -0.27 0.13 -0.53 0.67 1000
twi 1.20 0.80 0.40 -0.13 -1.47 -0.80 -0.80 -0.13 5000
wol -0.27 -0.13 0.93 0.93 0.13 0.80 0.40 0.27 1000
wol -0.53 -0.40 -0.40 -0.40 -0.40 -0.27 0.27 0.67 5000
wol 0.40 0.67 -0.67 -0.53 0.93 -0.40 0.93 0.00 10000
wol 0.40 0.67 -0.67 -0.53 0.93 -0.40 0.93 0.00 25100
xho -0.27 -1.20 0.27 0.93 0.67 0.00 1.07 1.87 1000
xho 0.13 -0.93 -0.40 -0.53 -1.73 0.27 -1.60 0.40 5000
xho -0.67 -0.27 -0.13 -0.67 -0.40 -0.13 -0.40 -1.07 10000
xho -0.67 -0.27 -0.13 -0.67 -0.53 0.40 0.00 -1.33 25100
yor -1.07 -0.67 -0.40 0.27 0.00 0.13 1.20 0.13 1000
yor 0.27 1.33 0.00 -0.13 0.80 -0.27 1.07 0.00 5000
yor -0.93 0.27 -0.27 -0.93 -1.20 -0.93 0.67 -0.93 10000
yor -0.93 0.27 -0.27 -0.93 -0.13 0.53 -0.40 0.27 25100
zul 0.27 -1.07 0.80 -0.80 -0.67 1.20 0.40 0.13 1000
zul 0.80 -0.67 -0.27 -0.80 0.67 0.67 0.27 -1.07 5000
zul -1.07 0.27 -0.93 -1.20 -0.53 0.40 0.00 -1.33 10000
zul -1.07 0.27 -0.93 -1.20 -0.93 0.67 0.27 -1.33 25100
Table 4: Full Native Robustness (Numeric Match $\Delta$). We report $\Delta_{\mathrm{NM}}=\mathrm{NM}_{\text{localized}}-\mathrm{NM}_{\text{translated}}$ across all sampled data fine-tunings for translated data ($x_{\text{trans}}$), English entity–replaced data ($x_{\text{ent}}$), auto-localized data ($\hat{x}_{\text{loc}}$), and all data combined. Positive values indicate higher robustness on localized benchmarks, negative values indicate stronger performance on English-centric benchmarks, and zero denotes no change.
Second, sample size plays a crucial role. For low-resource settings (e.g., 1k samples), both models display instability, with performance swings that highlight the difficulty of robustly adapting to local contexts with very limited data. As the number of samples increases (e.g., 25k), more stable trends appear, although the direction of improvement still varies by language. For instance, some languages (such as Kinyarwanda and Shona) show strong positive gains under localization, whereas others (such as Hausa and Luganda) exhibit persistent negative or mixed trends, indicating that language-specific features or annotation artifacts may influence results.
The “all data” setting, combining translated, entity-replaced, and localized data, yields more balanced robustness across languages. This indicates that hybrid augmentation can reduce model brittleness, though closing language gaps requires attention to data quality and localization type, not just scale.
These results underscore the interplay between training data, sampling size, and language characteristics, emphasizing the need to evaluate models on culturally grounded datasets rather than translations alone.
Appendix B Adopted Evaluation Prompts
We adapted three prompts from Adelani et al. (2024) and customized them to ensure that they return only numeric answers.
Prompt: Adopted prompt for evaluations prompt_1: "Question: {{lang}}
Return the number answer only. Do not provide an explanation.
Number Answer:"
prompt_2: "Give direct numerical answers for the question provided.
Question: {{lang}}
Do not provide an explanation.
Numeric Answer:"
prompt_3: "Solve the following math question.
Question: {{lang}}
Do not provide an explanation.
Numeric Answer:"
Appendix C Manual Annotation
For the human annotation shown in Table 2, annotators were asked to examine $x_{\text{ent}}$ and determine whether English-centric names, currencies, or organization names had been replaced. If such entities were replaced, the sample was labeled as Culturally Localized; if they remained, it was labeled as Not Localized. Since all languages share the same $x_{\text{ent}}$, the manual annotation was carried out once and applied across all languages.
Appendix D Prompt Templates for the Localization Pipeline
@@
Prompt: Entity Localization Task You are an expert linguistic assistant. Your task is to edit a sentence in a native language to match a change made in its English parallel. Here is the context: Original English: The original sentence. Original Native: The original translation of the English sentence in {native_lang}. Modified English: The English sentence has been edited. One or more words have been replaced. Your goal is to produce a Modified Native sentence by applying the exact same replacement to the Original Native sentence. Crucial Instructions: DO NOT re-translate the entire sentence. Only replace the specific words that were changed in the English version. Preserve the original grammar and structure of the native sentence as much as possible. Ensure the final Modified Native sentence is natural and grammatically correct in {native_lang}. Respond with ONLY the Modified Native sentence and nothing else. Example: Original English: Janet’s ducks lay 16 eggs per day. Original Native (French): Les canards de Janet pondent 16 œufs par jour. Modified English: Andrea’s ducks lay 16 eggs per day. Modified Native (French): Les canards d’Andrea pondent 16 œufs par jour. Your Task: Original English: {original_eng} Original Native ({native_lang}): {original_native} Modified English: {modified_eng} Modified Native ({native_lang}):
@@
Prompt: Direct Localization using LLMs without our pipeline You are an expert linguistic assistant specializing in localization. Your task is to localize the following English sentence into {target_lang_name}. Localization means adapting the text to the target culture and context by: Identifying and replacing English-specific entities (such as person names, organization names, and currency symbols/names) with culturally appropriate equivalents in {target_lang_name}. Ensuring the localized text sounds natural and grammatically correct in {target_lang_name}. Crucially, the localized text should be generated directly from the English sentence, without relying on any pre-existing native translation. English Sentence to Localize: {original_eng} Your Goal: Produce a Localized {target_lang_name} sentence. Instructions: Identify person names, organization names, and currency symbols/names in the English sentence. Replace these entities with culturally appropriate {target_lang_name} alternatives. If an entity does not have a direct cultural equivalent or is already culturally neutral, it may remain unchanged. Respond with ONLY the Localized {target_lang_name} sentence and nothing else. Localized {target_lang_name}:
Appendix E Presence of Cultural entities in our Training data
@@
Figure 5 illustrates the distribution of culturally relevant items within the 1,500 data samples selected for each language. Our analysis reveals that a considerable proportion of the translated items could not be localized to the target language, primarily because they did not contain the types of entities, such as personal names, organizations, or currencies, that form the basis of our localization strategy.
This proportion also helps explain why localized datasets did not improve all languages equally. For some languages, high-quality translations may already omit most culturally salient entities, limiting the added value of localization. In contrast, for languages such as Kinyarwanda, where out of 1500 translations only about 150 lacked cultural references despite high overall translation scores, localization introduced substantial additional signal. This variation across languages highlights the interaction between translation quality, cultural entity coverage, and the benefits of localization.
@@
Figure 5: Number of data samples without cultural entity replacements (out of 1500 selected), grouped by language in the training dataset.
Appendix F Translation quality - SSA-COMMET score
We leveraged the GSM8K dataset from Cobbe et al. (2021) and applied our pipeline to generate translated and localized versions. For translation, we used the NLLB-200-3.3B model NLLB Team et al. (2022). To ensure quality, we filtered out translations with an SSA-COMMET score below 0.65, a threshold we determined through manual analysis as providing a good balance between quality and coverage. While SSA-COMMET scores give a reliable indication of translation performance, we acknowledge that occasional errors in the text may remain.
@@ Appendix F Translation quality - SSA-COMMET score
Appendix G Error Analysis on RQ2: Native Language Outputs
Table 5 shows comparison of numeric answers versus non-numeric answers across prompts for different models. Because of this issue in the results making fair comparison challenging, we reported the best result across prompts instead of averaging them for Figure 3, similar to prior work kamruzzaman2025anger. The results are for 17 languages, 3 prompts, 3 experiments and 250 data each.
LLaMA-3-70B-Instruct
Numeric Non Numeric
prompt_1 12741 9
prompt_2 12369 381
prompt_3 11609 1141
GPT-4o-mini
Numeric Non Numeric
prompt_1 12734 16
prompt_2 12725 25
prompt_3 12670 80
Aya-expanse-32b
Numeric Non Numeric
prompt_1 12708 42
prompt_2 5469 7281
prompt_3 12286 464
Gemma-2-27b-it
Numeric Non Numeric
prompt_1 12750 0
prompt_2 12737 13
prompt_3 12734 16
Table 5: Error analysis on RQ2 (Native language outputs) of selected models.
Appendix H Code and Reproducibility
@@
Due to the challenges of obtaining numeric answers in low-resource mathematical evaluations, we adopted a strategy of extracting the best answer option and reporting numeric match scores based on the best-performing prompt for each model (Table 3). In contrast, since we were able to obtain a sufficient number of numeric answers for English math word problems (MWPs), we averaged the results across three prompts and report these in Table 4.
Section Parameters
Model model_name_or_path: [google/gemma-2-9b-it, meta-llama/Meta-Llama-3-8B-Instruct]
# choose one
trust_remote_code: true
Method stage: sft
do_train: true
finetuning_type: lora
lora_rank: 8
lora_target: all
Dataset dataset: [gsm8k-math-localized, gsm8k-math-english, gsm8k-math-models-translated]
# choose one or all
template: [gemma2, llama3]
# choose one
cutoff_len: 2048
max_samples: #1000,5000,10000,none
overwrite_cache: true
preprocessing_num_workers: 16
dataloader_num_workers: 4
Output output_dir: #some directory
logging_steps: 10
save_steps: 500
plot_loss: true
overwrite_output_dir: true
save_only_model: false
report_to: none
Train per_device_train_batch_size: 1
gradient_accumulation_steps: 8
learning_rate: 1.0e-4
num_train_epochs: 3.0
lr_scheduler_type: cosine
warmup_ratio: 0.1
bf16: true
ddp_timeout: 180000000
resume_from_checkpoint: null
Table 6: Configuration for fine-tuning gemma-2-9b-it and Meta-Llama-3-8B-Instruct